OpenStudy (anonymous):

Gina, Sam and Robby all rented movies from the same video store. They each rented some dramas, comedies and documentaries. Gina rented 11 movies total. Sam rented twice as many dramas, three times as many comedies, and twice as many documentaries as Gina. He rented 27 movies total. If Robby rented 19 movies total with the same number of dramas, twice as many comedies, and twice as many documentaries as Gina, how many movies of each type did Gina rent?

OpenStudy (anonymous):

A.) 3 dramas, 5 comedies and 3 documentaries B.) 1 dramas, 4 comedies and 6 documentaries C.) 2 dramas, 6 comedies and 3 documentaries D.) 4 dramas, 3 comedies and 4 documentaries

OpenStudy (sriram):

A

OpenStudy (anonymous):

Question 1] let d = number of dramas Gina rented, c = # of comedies, and t = # of documentaries: [equation a] d + c + t = 11 [equation b] 2d + 3c + 2t = 27 [equation c] d + 2c + 2t = 19 if we subtract [a] from [c] we get: [equation d] c + t = 8 if we subtract [c]*2 from [b] we get: [equation e] -c - 2t = -11 or [equation e] c + 2t = 11 Now we can subtract [e] from [d] to solve for t: t = 3 Now put in 3 for t in [d] or [e] to solve for c: c + 2(3) = 11 c + 6 = 11 c = 5 So she rented 5 comedies and 3 documentaries which leaves 3 dramas (which is answer A)